# Pinning control of chimera states in systems with higher-order interactions

### v1

@@ Abstract
Understanding and controlling the mechanisms behind synchronization phenomena is of paramount importance in nonlinear science. In particular, chimera states, patterns in which order and disorder coexist simultaneously, continue to puzzle scholars, due to their elusive nature. Recently, it has been shown that higher-order interactions greatly promote the onset of chimera states, which are easier to found and more resilient when the system units interact in group. In this work, we show that the higher-order framework is fertile not only for the emergence of chimera states, but also for their control. Via pinning control, a technique consisting in applying a forcing to a subset of the nodes, we are able to trigger the emergence of chimera states with only a small fraction of controlled nodes, at striking contrast with the case without higher-order interactions. We show that our setting is robust for different higher-order topologies and types of pinning control and, finally, we give a heuristic interpretation of the results via phase reduction theory. Our numerical and theoretical results provide further understanding on how higher-order interactions shape nonlinear dynamics.
@@ I Introduction
Understanding the mechanisms underlying self-organization phenomena on networks is a paramount task in the study of complex systems, which is complemented by the development of efficient methods to control such dynamics D’Souza et al. (2023). This is particularly relevant in the framework of synchronization dynamics, where, depending on the applications, it is fundamental to achieve a synchronized state, e.g., in power grids Sajadi et al. (2022), or to break it into an asynchronous one, e.g., in neuroscience Asllani et al. (2018), where synchronization is often associated to pathological states. The network framework remains still relevant in the modeling of complex systems, nonetheless, over the past years scholars have started considering more complex structures such as hypergraphs and simplicial complexes Battiston et al. (2020, 2021); Bianconi (2021); Boccaletti et al. (2023); Bick et al. (2023a). This is because networks do not capture interactions beyond the pairwise setting, i.e., two-by-two, while many systems have shown evidence of higher-order, i.e., group, interactions Battiston et al. (2020, 2021). Examples come from, but are not limited to, neuroscience Andrew et al. (2021); Wang et al. (2013); Petri et al. (2014); Sizemore et al. (2018), ecology Grilli et al. (2017); Iacopini et al. (2024) and social behaviors Centola et al. (2018). Higher-order interactions have been proven to greatly affect the collective behavior, for instance, in random walks Carletti et al. (2020a); Schaub et al. (2020), synchronization dynamics Tanaka and Aoyagi (2011); Skardal and Arenas (2020); Millán et al. (2020), contagion Iacopini et al. (2019) and pattern formation Carletti et al. (2020b); Muolo et al. (2023a), to name just a few. Given the ubiquity of group interactions Battiston et al. (2020); Bianconi (2021); Battiston et al. (2021); Boccaletti et al. (2023); Bick et al. (2023a), it is important to understand how to control the dynamics in such systems. While significant progress has been made in the control of networks Liu et al. (2011); Liu and Barabási (2016), the investigation into the control of systems with higher-order interactions has only recently begun Chen et al. (2021); De Lellis et al. (2022, 2023); Xia and Xiang (2024).
@@
The focus of this work is an intriguing type of synchronization pattern called chimera state, which consists of the coexistence of coherent and incoherent domains of oscillations. Coexistence of coherence and incoherence was first observed by Kaneko for globally coupled chaotic maps Kaneko (1990) and was then found in several numerical settings with global Hakim and Rappel (1992); Nakagawa and Kuramoto (1993); Chabanol et al. (1997) and nonlocal Kuramoto (1995); Kuramoto and Nakao (1996, 1997); Kuramoto et al. (1998, 2000) coupling schemes. Despite all the previous research on the subject, the work that historically is considered to be the first to characterize the emergence of chimera states is the well-known paper by Kuramoto and Battogtokh Kuramoto and Battogtokh (2002), made popular by a successive work of Abrams and Strogatz, who, with a creative intuition, compared the coexistence of different dynamical state to the chimera, a mythological creature in which parts of different animals coexisted Abrams and Strogatz (2004). Besides the pure theoretical relevance of such an astonishing phenomenon, a great part of the interest has been generated by the existence of analogous patterns in real systems: for instance, in Josephson junctions Domínguez and Cerdeira (1993) and electronic circuits Gambuzza et al. (2014, 2020), laser Hagerstrom et al. (2012), mechanical Martens et al. (2013) and nano-electromechanical systems Matheny et al. (2019), to name a few. Particular attention has been devoted to neuroscience Chouzouris et al. (2018); Majhi et al. (2019), specifically to unihemispheric sleeping patterns in animals Rattenborg et al. (2000). Except for some particular configurations in which robust chimera patterns are induced by the network structure Asllani et al. (2022); Muolo et al. (2023b), in both numerical and experimental settings chimera states are often elusive and characterized by a rather short lifetime. Hence, there is a vast literature on networked systems, consisting in looking for different settings (e.g., parameter ranges, network topologies, coupling configurations, etc.) making such patterns easier to find and with a longer lifetime. Moreover, after the first definition by Kuramoto and Battogtokh Kuramoto and Battogtokh (2002), several kinds of chimera states have been defined, e.g., amplitude chimeras Zakharova et al. (2016) or phase chimeras Zajdela and Abrams (2023). We will not thoroughly discuss such studies, inviting the interested reader to consult a book Zakharova (2020) and a review Parastesh et al. (2021) on the subject. In the context of higher-order interactions, chimera states have been proven to be enhanced in some pioneering works considering both pairwise and higher-order interactions Kundu and Ghosh (2022a); Li et al. (2023); Bick et al. (2023b). This claim was further corroborated in Muolo et al. (2024) for systems with pure higher-order interactions, where the emergence of chimera states on higher-order topologies is compared with the absence of such patterns when the interactions are pairwise.
@@
In this work, we consider the setting studied in Muolo et al. (2024) and implement a control to further trigger the emergence of chimera states. Our control approach will rely on the so called pinning control, a technique used to drive networks onto a desired dynamical state by using a control input applied to a small subset of nodes Grigoriev et al. (1997); Sorrentino et al. (2007). Such technique has been successfully used in the framework of opinion dynamics Lo Iudice et al. (2022); Ancona et al. (2023), epidemics Du Toit and Craig (2015); Yang et al. (2019), pattern formation Buscarino et al. (2019) and synchronization dynamics Porfiri and di Bernardo (2008); Yu et al. (2013); Gambuzza and Frasca (2016), to name a few. The latter includes the control of chimera states with pairwise interactions, which we hereby extend to the higher-order framework. Indeed, in Gambuzza and Frasca (2016) it was shown that it is possible to trigger the emergence of chimera states via pinning. Nonetheless, to achieve such task, at least half of the network nodes need to be controlled. In what follows, after introducing the setting in Sec. II, we show that higher-order interactions considerably facilitate the work of the controllers and chimera states can be obtained by controlling only a small fraction of the nodes. Such results are shown in Sec. III for two different kinds of pinning approaches, that we named additive pinning and parametric pinning. Moreover, we show that, rather than the number of nodes, what matters is the size of the hyperedges, i.e., the group of nodes interacting with each other. Then, before the discussion of some potential future directions, we give a heuristic interpretation of the results based on the phase reduction theory Nakao (2016) in Sec. IV.
@@ II The model and the setting
In this Section, we introduce the system exhibiting chimera states, which is analogous to that studied in Muolo et al. (2024). We consider coupled Stuart-Landau oscillators, a paradigmatic model for the study of synchronization dynamics, given that it is the normal form of the supercritical Hopf-Andronov bifurcation Nakao (2014). The coupling takes place through pure higher-order interactions, namely, by mean of a higher-order topology called nonlocal hyperring, which is a generalization of the nonlocal pairwise coupling Muolo et al. (2024). The type of chimera state that we will hereby consider is that of phase chimeras, states have been first observed by Zajdela and Abrams Zajdela and Abrams (2023), which consist in oscillation patterns where the amplitude and the frequency of each oscillator are the same, but the phases exhibit a chimera behavior, i.e., a part of the oscillators have the same phase, while the other phases are distributed along the unit circle. The peculiarity of such pattern is that, once obtained, it does not vary, because the frequency is the same for all the oscillators. Hence, we would observe the same exact pattern after each period. For this reason, we find the description given by Zajdela and Abrams, ”frozen patterns of disorder”, perfectly fitting. The reader can find a thorough characterization of these patterns in the Refs. Zajdela and Abrams (2023); Muolo et al. (2024). On a side, let us note that multitailed phase chimeras have only been found in the pairwise setting Zajdela and Abrams (2023), but not yet in the higher-order one. In what follows, every chimera state discussed and shown in the figures will be a phase chimera. For sake of simplicity, we will refrain from using the word ”phase” and will call them simply ”chimeras”.
@@ II.1 Stuart-Landau oscillators coupled via nonlocal hyperrings
where $\alpha$ is a bifurcation parameter and $\omega$ is the frequency of the oscillators. Let us stress that the units are homogeneous, meaning that the parameters $\alpha$ and $\omega$ are the same for each and every system. Each isolated system exhibits a stable limit cycle for $\alpha>0$, which is the case we will consider throughout this study.
@@
To model the higher-order interactions we use a generalization of the links (or edges) called hyperedges, whose structure can be encoded by using adjacency tensors, a generalization of the adjacency matrices for networks Battiston et al. (2020). From the literature dealing with simplicial complexes Bianconi (2021), we adopt the convention that a hyperedge of $(d+1)$ nodes (i.e, encoding a $(d+1)$-body interaction) is called a $d$-hyperedge. As an example, let us consider the $3$-rd order [^1] Let us note that in tensor algebra the order (or rank) of a tensor is given by its indices, e.g., a scalar is a 00-rank tensor, a vector is a $1$-rank tensor, a matrix is a $2$-rank tensor, etc. Hence, $A^{(3)}=\{A_{ijkl}^{(3)}\}$ would be a $4$-rank tensor and the adjacency matrix, given that it is a matrix, a $2$-rank tensor. However, here we follow that notation that is most commonly used in the literature on higher-order interactions. adjacency tensor (i.e., encoding $4$-body interactions) $A^{(3)}=\{A_{ijkl}^{(3)}\}$. We have that $A_{ijkl}^{(3)}=1$ if nodes $i,j,k,l$ are part of the same hyperedge and 00 otherwise. This is the analogous of the adjacency matrix for networks, which is, indeed, a $1$-st order adjacency tensor. The chosen higher-order topology is that of nonlocal $d$-hyperrings, a pure higher-order structure introduced in Muolo et al. (2024) and consisting of hyperedges of size $d+1$ ($d$ is the order of the interaction) attached through one node and set in a circular structure. In Fig. 1, we depict a $3$-hyperring, in panel a), together with its pairwise counterpart, namely, the clique-projected network (cpn) obtained by projecting each hyperedge into a clique having the same size, in panel b). Observe that a hyperring is a uniform hypergraph, the hyperedges all having the same size.
@@
Following the analyses carried out in previous works Gambuzza et al. (2020); Muolo et al. (2024), we assume the coupling to involve only the first state variable of each oscillator, i.e., $x$. Let us stress that other coupling configurations can be considered, as discussed in Appendix A. With the above assumption, the equations for systems (1) coupled via a generic $d$-hyperring read
$\begin{cases}\displaystyle\dot{x}_{i}=f(x_{i},y_{i})+\varepsilon\sum_{j_{1},...,j_{d}=1}^{N}A_{i,j_{1},...,j_{d}}^{(d)}\Big{(}h^{(d)}(x_{j_{1}},...,x_{j_{d}})-h^{(d)}(x_{i},...,x_{i})\Big{)},\\ \\ \displaystyle\dot{y}_{i}=g(x_{i},y_{i})+\varepsilon\sum_{j_{1},...,j_{d}=1}^{N}A_{i,j_{1},...,j_{d}}^{(d)}\Big{(}h^{(d)}(x_{j_{1}},...,x_{j_{d}})-h^{(d)}(x_{i},...,x_{i})\Big{)},\end{cases}$ (2)
@@
where $\varepsilon>0$ is the coupling strength [^2] Note that such configuration involves only interactions of order $d$, i.e., $(d+1)$-body, hence it is not necessary to denote the coupling strength with $\varepsilon_{d}$.. Because we consider identical oscillators, together with the fact that the coupling is diffusive like, i.e., it vanishes when all oscillators are in the same state, system (2) admits a fully synchronous solution. Such a coupling is a special type of non-invasive interaction Gambuzza et al. (2021). Moreover, we will consider coupling functions such that the higher-order interaction cannot be decomposed into pairwise ones [^3] It was shown by Neuhäuser et al. that the higher-order coupling functions need to be nonlinear, otherwise the many-body interaction can be decomposed into pairwise ones Neuhäuser et al. (2020). Successively, it was further shown that such assumption may not be enough and, if the nonlinear functions $h$ are the sum of nonlinear terms which separately account for each unit, e.g., $h^{(d)}(x_{j_{1}},...,x_{j_{d}})=h(x_{j_{1}})+...+h(x_{j_{d}})$, they can still be decomposed into pairwise ones Muolo et al. (2023a).. Eq. (2) can be rewritten in compact form for each unit $i$ as
$\dot{\vec{X}}_{i}=\vec{F}(\vec{X}_{i})+\mathbb{D}\sum_{j_{1},...,j_{d}=1}^{N}A_{i,j_{1},...,j_{d}}^{(d)}\Big{(}\vec{H}^{(d)}(\vec{X}_{j_{1}},...,\vec{X}_{j_{d}})-\vec{H}^{(d)}(\vec{X}_{i},...,\vec{X}_{i})\Big{)},$ (3)
@@
where $\vec{X}_{i}=(x_{i},y_{i})^{\top}$, $\vec{F}=(f,g)^{\top}$, $\vec{H}^{(d)}=(h^{(d)},h^{(d)})^{\top}$ and $\mathbb{D}=\varepsilon D=\varepsilon\begin{bmatrix}1&0\\ 1&0\end{bmatrix}$. As stated previously, in Appendix A we report additional results for different coupling matrices $D$.
@@
To highlight the effects of higher-order interactions, we will perform a comparison between the dynamics on the hyperring and on its respective clique-projected network (cpn), as in Muolo et al. (2024). The equations for the dynamics with pairwise interactions are
$\begin{cases}\displaystyle\dot{x}_{i}=f(x_{i},y_{i})+\varepsilon\sum_{j=1}^{N}A_{i,j}^{(1)}\Big{(}h^{cpn}(x_{j})-h^{cpn}(x_{i})\Big{)},\\ \\ \displaystyle\dot{y}_{i}=g(x_{i},y_{i})+\varepsilon\sum_{j=1}^{N}A_{i,j}^{(1)}\Big{(}h^{cpn}(x_{j})-h^{cpn}(x_{i})\Big{)},\end{cases}$ (4)
@@
We will perform pinning control on hyperrings of $4$ different orders, namely, $3$-,$4$-,$5$- and $6$-hyperrings, involving $4$-,$5$-,$6$- and $7$-body interactions, respectively. In line with other previous works, e.g., Gambuzza et al. (2021), we will consider odd-degree polynomials as coupling functions. The coupling functions for each hyperring and its clique-projected network are the following
@@
where the adjacency tensor accounts for all the permutations of the indexes. As an example, let us explicit the equations for the $4$-body case ($3$-hyperring) and its corresponding clique-projected network. The equations for a $3$-hyperring are the following
$\begin{cases}\displaystyle\dot{x}_{i}=\alpha x_{i}-\omega y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)x_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{i,j_{1},j_{2},j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)},\\ \\ \displaystyle\dot{y}_{i}=\omega x_{i}+\alpha y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)y_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{i,j_{1},j_{2},j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)},\end{cases}$ (6)

### v2

@@ Abstract
Understanding and controlling the mechanisms behind synchronization phenomena is of paramount importance in nonlinear science. In particular, the emergence of chimera states, patterns in which order and disorder coexist simultaneously, continues to puzzle scholars, due to its elusive nature. Recently, it has been shown that higher-order (many-body) interactions greatly enhance the presence of chimera states, which are easier to be found and more persistent. In this work, we show that the higher-order framework is fertile not only for the emergence of chimera states, but also for its control. Via pinning control, a technique consisting in applying a forcing to a subset of the nodes, we are able to trigger the emergence of chimera states with only a small fraction of controlled nodes, at striking contrast with the case without higher-order interactions. We show that our setting is robust for different higher-order topologies and types of pinning control and, finally, we give a heuristic interpretation of the results via phase reduction theory. Our numerical and theoretical results provide further understanding on how higher-order interactions shape collective behaviors in nonlinear dynamics.
@@ I Introduction
Understanding the mechanisms underlying self-organization phenomena on networks is a paramount task in the study of complex systems, which is complemented by the development of efficient methods to control such dynamics d2023controlling . This is particularly relevant in the framework of synchronization dynamics, where, depending on the applications, it is fundamental to achieve a synchronized state, e.g., in power grids sajadi2022synchronization , or to break it into an asynchronous one, e.g., in neuroscience asllani2018minimally , where synchronization is often associated to pathological states. The network framework remains still relevant in the modeling of complex systems, nonetheless, over the past years scholars have started considering more complex structures such as hypergraphs and simplicial complexes battiston2020networks ; battiston2021physics ; bianconi2021higher ; boccaletti2023structure ; bick2023higher ; millan2025topology . This is because networks do not capture interactions beyond the pairwise setting, i.e., two-by-two, while many systems have shown evidence of higher-order, i.e., group, interactions battiston2020networks ; battiston2021physics . Examples come from, but are not limited to, neuroscience rosen1989 ; wang2013 ; petri2014homological ; sizemore2018cliques , ecology grilli_allesina ; iacopini2024not and social behaviors centola2018experimental . Higher-order interactions have been proven to greatly affect the collective behavior, for instance, in random walks carletti2020random ; schaub2020random , synchronization dynamics tanaka2011multistable ; skardal2019higher ; millan2020explosive , contagion iacopini2019simplicial and pattern formation carletti2020dynamical ; muolo2024turing , to name just a few. Given the ubiquity of group interactions battiston2020networks ; battiston2021physics ; bianconi2021higher ; boccaletti2023structure ; bick2023higher , it is important to understand how to control the dynamics in such systems. While significant progress has been made in the control of networks liu2011controllability ; liu2016control , the investigation into the control of systems with higher-order interactions has only recently begun chen2021controllability ; de2022pinning ; de2023pinning ; xia2024pinning ; moriame2025hamiltonian .
@@
The focus of this work is an intriguing type of synchronization pattern called chimera state, which consists of the coexistence of coherent and incoherent domains of oscillations. Coexistence of coherence and incoherence was first observed by Kaneko for globally coupled chaotic maps kaneko and was then found in several numerical settings with global hakim1992dynamics ; nakagawa1993collective ; chabanol1997collective and nonlocal kuramoto1995scaling ; kuramoto1996origin ; kuramoto1997power ; kuramoto1998multiaffine ; kuramoto2000multi coupling schemes. Despite all the previous research on the subject, the work that historically is considered to be the first to characterize the emergence of chimera states is the well-known paper by Kuramoto and Battogtokh kuramoto2002coexistence , made popular by a successive work of Abrams and Strogatz, who, with a creative intuition, compared the coexistence of different dynamical state to the chimera, a mythological creature in which parts of different animals coexisted abrams2004chimera . Besides the pure theoretical relevance of such an astonishing phenomenon, a great part of the interest has been generated by the existence of analogous patterns in real systems: for instance, in Josephson junctions cerdeira_prl and electronic circuits vale_chimeras ; gambuzza2020experimental , laser hagerstrom2012experimental , mechanical martens2013chimera and nano-electromechanical systems matheny2019exotic , to name a few. Particular attention has been devoted to neuroscience chimera_neuro ; majhi2019chimera , specifically to unihemispheric sleeping patterns in animals rattenborg2000behavioral . Except for some particular configurations in which robust chimera patterns are induced by the network structure bram_malb_chim ; muolo_2023_chimera ; asllani2025pattern , and for which a rigorous stability analysis can be carried out, in both numerical and experimental settings chimera states are often elusive and characterized by a rather short lifetime. Note that, if the coupling is global (all-to-all), chimeras may not always be transient in the thermodynamic limit. Hence, there is a vast literature on networked systems, consisting in looking for different settings (e.g., parameter ranges, network topologies, coupling configurations, etc.) making such patterns easier to find and with a longer lifetime. Moreover, after the first definition by Kuramoto and Battogtokh kuramoto2002coexistence , several kinds of chimera states have been defined, e.g., amplitude chimeras zakharova2016amplitude or phase chimeras zajdela_abrams . We will not thoroughly discuss such studies, inviting the interested reader to consult a book zakharova2020chimera and a review parastesh2021chimeras on the subject. In the context of higher-order interactions, chimera states have been proven to be enhanced in some pioneering works considering both pairwise and higher-order interactions kundu2022higher ; ghosh_chimera2 ; bick_nonlocal1 . This claim was further corroborated in muolo2024phase for systems with pure higher-order interactions, where the emergence of chimera states on higher-order topologies was compared with the absence of such patterns when the interactions are pairwise.
@@
Several techniques have been developed to control the emergence of chimera states on networks, including, but not limited to, gradient dynamics bick2015controlling , feedback control omelchenko2018optimal ; omelchenko2019control , delay semenov2016deterministic ; ghosh2018engineering ; ghosh2019taming , pace makers ruzzene2020remote , and time varying networks majhi2022oscillation , to name a few. In this work, we consider the setting studied in muolo2024phase and implement a control to further trigger the emergence of chimera states. Our control approach will rely on the so called pinning control, a technique used to drive networks onto a desired dynamical state by using a control input applied to a small subset of nodes grigoriev1997pinning ; sorrentino2007controllability . Such technique has been successfully used in the framework of opinion dynamics iudice2022bounded ; ancona2023influencing , epidemics du2015selective ; yang2019feedback , pattern formation buscarino2019turing and synchronization dynamics porfiri2008criteria ; yu2013synchronization ; gambuzza2016pinning , to name a few. The latter includes the control of chimera states with pairwise interactions, which we hereby extend to the higher-order framework. Indeed, in gambuzza2016pinning it was shown that it is possible to trigger the emergence of chimera states via pinning. Nonetheless, to achieve such task, at least half of the network nodes need to be controlled. In what follows, after introducing the setting in Sec. II, we show that higher-order interactions considerably facilitate the work of the controllers and chimera states can be obtained by controlling only a small fraction of the nodes. Such results are shown in Sec. III for two different kinds of pinning approaches, that we named additive pinning and parametric pinning. Moreover, we show that, rather than the number of nodes, what matters is the size of the hyperedges, i.e., the group of nodes interacting with each other. Then, before the discussion of some potential future directions, we give a heuristic interpretation of the results based on the theory of phase reduction nakao2016phase in Sec. IV.
@@ II The model and the setting
In this Section, we introduce the system exhibiting chimera states, which is analogous to that studied in muolo2024phase . We consider coupled Stuart-Landau oscillators, a paradigmatic model for the study of synchronization dynamics, given that it is the normal form of the supercritical Hopf-Andronov bifurcation nakao2014complex . The coupling takes place through pure higher-order interactions, namely, by mean of a higher-order topology called nonlocal hyperring, which is a generalization of the nonlocal pairwise coupling muolo2024phase . The type of chimera state that we will hereby consider is that of phase chimeras, states have been first observed by Zajdela and Abrams zajdela_abrams , which consist in oscillation patterns where the amplitude and the frequency of each oscillator are the same, but the phases exhibit a chimera behavior, i.e., a part of the oscillators have the same phase, while the other phases are distributed along the unit circle. The peculiarity of such pattern is that, once obtained, it does not vary, because the frequency is the same for all the oscillators. Hence, we would observe the same exact pattern after each period. For this reason, we find the description given by Zajdela and Abrams, ”frozen patterns of disorder”, perfectly fitting. The reader can find a thorough characterization of these patterns in the Refs. zajdela_abrams ; muolo2024phase , which are not completely ”frozen” in the latter, the coupling not being all-to-all. On a side, let us note that multitailed phase chimeras have only been found in the pairwise setting zajdela_abrams , but not yet in the higher-order one. In what follows, every chimera state discussed and shown in the figures will be a phase chimera. For sake of simplicity, we will refrain from using the word ”phase” and will call them simply ”chimeras”.
@@ II.1 Stuart-Landau oscillators coupled via nonlocal hyperrings
where $\alpha$ is a bifurcation parameter and $\omega$ is the frequency of the oscillators. Let us stress that the units are homogeneous, meaning that the parameters $\alpha$ and $\omega$ are the same for each and every system. Each isolated system exhibits a stable limit cycle for $\alpha>0$, which is the case we will consider throughout this study. Note that the above formulation of the Stuart-Landau equations comes from the Complex Ginzburg-Landau Equation (CGLE), from which real and imaginary parts are taken, which are variables $x$ and $y$, respectively nakao2014complex .
@@
To model the higher-order interactions we use a generalization of the links (or edges) called hyperedges, whose structure can be encoded by using adjacency tensors, a generalization of the adjacency matrices for networks battiston2020networks . From the literature dealing with simplicial complexes bianconi2021higher , we adopt the convention that a hyperedge of $(d+1)$ nodes (i.e, encoding a $(d+1)$-body interaction) is called a $d$-hyperedge. As an example, let us consider the $3$-rd order [^1] Let us note that in tensor algebra the order (or rank) of a tensor is given by its indices, e.g., a scalar is a 00-rank tensor, a vector is a $1$-rank tensor, a matrix is a $2$-rank tensor, etc. Hence, $A^{(3)}=\{A_{ijkl}^{(3)}\}$ would be a $4$-rank tensor and the adjacency matrix, given that it is a matrix, a $2$-rank tensor. However, here we follow that notation that is most commonly used in the literature on higher-order interactions. adjacency tensor (i.e., encoding $4$-body interactions) $A^{(3)}=\{A_{ijkl}^{(3)}\}$. We have that $A_{ijkl}^{(3)}=1$ if nodes $i,j,k,l$ are part of the same hyperedge and 00 otherwise. This is the analogous of the adjacency matrix for networks, which is, indeed, a $1$-st order adjacency tensor. The chosen higher-order topology is that of nonlocal $d$-hyperrings, a pure higher-order structure introduced in muolo2024phase and consisting of hyperedges of size $d+1$ ($d$ is the order of the interaction) attached through one node and set in a circular structure. In Fig. 1, we depict a $3$-hyperring, in panel a), together with its pairwise counterpart, namely, the clique-projected network (cpn) obtained by projecting each hyperedge into a clique having the same size, in panel b). Observe that a hyperring is a uniform hypergraph, the hyperedges all having the same size.
@@
Following the analyses carried out in previous works gambuzza2020experimental ; muolo2024phase , we assume the coupling to involve only the first state variable of each oscillator, i.e., $x$. Let us stress that other coupling configurations can be considered, as discussed in Appendix A. With the above assumption, the equations for systems (1) coupled via a generic $d$-hyperring read
$\begin{cases}\displaystyle\dot{x}_{i}=f(x_{i},y_{i})+\varepsilon\sum_{j_{1},...,j_{d}=1}^{N}A_{ij_{1}...j_{d}}^{(d)}\Big{(}h^{(d)}(x_{j_{1}},...,x_{j_{d}})-h^{(d)}(x_{i},...,x_{i})\Big{)},\\ \\ \displaystyle\dot{y}_{i}=g(x_{i},y_{i})+\varepsilon\sum_{j_{1},...,j_{d}=1}^{N}A_{ij_{1}...j_{d}}^{(d)}\Big{(}h^{(d)}(x_{j_{1}},...,x_{j_{d}})-h^{(d)}(x_{i},...,x_{i})\Big{)},\end{cases}$ (2)
@@
where $\varepsilon>0$ is the coupling strength [^2] Note that such configuration involves only interactions of order $d$, i.e., $(d+1)$-body, hence it is not necessary to denote the coupling strength with $\varepsilon_{d}$.. Because we consider identical oscillators, together with the fact that the coupling is diffusive-like, i.e., it vanishes when all oscillators are in the same state, system (2) admits a fully synchronous solution. Such a coupling is a special type of non-invasive interaction gambuzza2021stability . Moreover, we will consider coupling functions such that the higher-order interaction cannot be decomposed into pairwise ones [^3] It was shown by Neuhäuser et al. that the higher-order coupling functions need to be nonlinear, otherwise the many-body interaction can be decomposed into pairwise ones neuhauser2020multibody . Successively, it was further shown that such assumption may not be enough and, if the nonlinear functions $h$ are the sum of nonlinear terms which separately account for each unit, e.g., $h^{(d)}(x_{j_{1}},...,x_{j_{d}})=h(x_{j_{1}})+...+h(x_{j_{d}})$, they can still be decomposed into pairwise ones muolo2023turing .. Eq. (2) can be rewritten in compact form for each unit $i$ as
$\dot{\vec{X}}_{i}=\vec{F}(\vec{X}_{i})+\mathbb{D}\sum_{j_{1},...,j_{d}=1}^{N}A_{ij_{1}...j_{d}}^{(d)}\Big{(}\vec{H}^{(d)}(\vec{X}_{j_{1}},...,\vec{X}_{j_{d}})-\vec{H}^{(d)}(\vec{X}_{i},...,\vec{X}_{i})\Big{)},$ (3)
@@
where $\vec{X}_{i}=(x_{i},y_{i})^{\top}$, $\vec{F}=(f,g)^{\top}$, $\vec{H}^{(d)}=(h^{(d)},h^{(d)})^{\top}$ and $\mathbb{D}=\varepsilon D=\varepsilon\begin{bmatrix}1&0\\ 1&0\end{bmatrix}$. As previously stated, in Appendix A we report additional results for different coupling matrices $D$.
@@
To highlight the effects of higher-order interactions, we will perform a comparison between the dynamics on the hyperring and on its respective clique-projected network (cpn), as in muolo2024phase . The equations for the dynamics with pairwise interactions are
$\begin{cases}\displaystyle\dot{x}_{i}=f(x_{i},y_{i})+\varepsilon\sum_{j=1}^{N}A_{ij}^{(1)}\Big{(}h^{cpn}(x_{j})-h^{cpn}(x_{i})\Big{)},\\ \\ \displaystyle\dot{y}_{i}=g(x_{i},y_{i})+\varepsilon\sum_{j=1}^{N}A_{ij}^{(1)}\Big{(}h^{cpn}(x_{j})-h^{cpn}(x_{i})\Big{)},\end{cases}$ (4)
@@
We will perform pinning control on hyperrings of $4$ different orders, namely, $3$-,$4$-,$5$- and $6$-hyperrings, involving $4$-,$5$-,$6$- and $7$-body interactions, respectively. For what concerns the functional form of the nonlinear coupling, we will choose odd-degree polynomials. This is because when Stuart-Landau (SL) oscillators are coupled through higher-order couplings of polynomial type with even degree, they behave as if there were no coupling. Indeed, through phase reduction theory, it can be shown that for oscillators with symmetry properties, such as the SL, higher-order couplings of polynomial type with even degree do not have any effect on the phase model at the first order approximation leon2025theory . The coupling functions for each hyperring and its clique-projected network are the following
@@
where the adjacency tensor accounts for all the permutations of the indexes. Note that the pairwise coupling for the systems on the clique-projected networks of the $4$- and $5$-hyperrings have the same functional form. We will come back to this point when discussing the dynamics of the systems with pairwise interactions. As an example, let us now explicitly write the equations for the $4$-body case ($3$-hyperring) and its corresponding clique-projected network. The equations for a $3$-hyperring are the following
$\begin{cases}\displaystyle\dot{x}_{i}=\alpha x_{i}-\omega y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)x_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)},\\ \\ \displaystyle\dot{y}_{i}=\omega x_{i}+\alpha y_{i}-\left(x_{i}^{2}+y_{i}^{2}\right)y_{i}+\varepsilon\sum_{j_{1},j_{2},j_{3}}^{N}A_{ij_{1}j_{2}j_{3}}^{(3)}\Big{(}x_{j_{1}}x_{j_{2}}x_{j_{3}}-x_{i}^{3}\Big{)},\end{cases}$ (6)